OpenStudy (anonymous):

calc II: find the taylor series (at x=0) of cos(x^2) can anyone help me i need reassurance to make sure my answers is correct ?

OpenStudy (amistre64):

whatcha got so far?

OpenStudy (anonymous):

didn't do it yet im doing it now i just did the first problem which was cos(x)

OpenStudy (amistre64):

wouldnt x=0 be a Maclaurin series?

OpenStudy (anonymous):

can you explain in your way of understanding when i would use Maclaurin series and taylor and which would be easier because my professor said to find it using taylor series

OpenStudy (anonymous):

as well as the answer

OpenStudy (anonymous):

Maclaurin series is just s special case of taylor series when it's centered at 0.

OpenStudy (amistre64):

its been awhile, but i think when f(a), that taylor series is used; but when a=0; then f(0); and the Maclaurin is used...

OpenStudy (anonymous):

because ive got a final tomorrow and im freaking out been studying and idk everything that im supposed to yet also yes i think thats what he said

OpenStudy (amistre64):

cos(0) -sin(0)-cos(0)+sin(0)+cos(0)....like that rght?

OpenStudy (amistre64):

forgot the factorials....

OpenStudy (amistre64):

cos(x^2) = 1- sin(0)x -cos(0)x^2+sin(0)x^3+cos(0)x^4 -------- -------- --------... 2! 3! 4!

OpenStudy (amistre64):

right?

OpenStudy (amistre64):

every other one goes to zero since sin(0)=0... which leaves us with: 1 -1 +1 -1 +1 -1 +1 -- -- -- -- -- -- ... right? 2! 4! 6! 8! 10! 12!

OpenStudy (amistre64):

x^2...x^4....x^6....

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

i havent finished yet

OpenStudy (anonymous):

I think the point here is the derivatives. I have does up to 8th derivatives, all are zeros at x=0 except the fourth and the eighth.

OpenStudy (amistre64):

i got no idea what to do after that :)

OpenStudy (anonymous):

Oh I got an idea. We can derive the Taylor series of cos(x^2) from that of cosx. Since tge expansion of cosx is well known

OpenStudy (anonymous):

So the Taylor series of cos x centered at 0 is: \[\cos x=\sum_{n=0}^{\infty}(-1)^n{x^{2n} \over (2n)!}\] Then:\[\cos(x^2)=\sum_{n=0}^{\infty}(-1)^n{(x^2)^{2n} \over (2n)!}=\sum_{n=0}^{\infty}(-1)^n{x^{4n} \over (2n)!}\]

OpenStudy (anonymous):

Do you know how to find the Taylor series of cos x at x=0?

OpenStudy (anonymous):

yes i did alread

OpenStudy (anonymous):

here

OpenStudy (anonymous):

Good. So, I just used that expansion and replaced x by x^2.

OpenStudy (anonymous):

F^(1) (x^2)=-sin x^2?

OpenStudy (anonymous):

f2 = -cosx^2?

OpenStudy (anonymous):

What's that?

OpenStudy (anonymous):

derivatives of the function

OpenStudy (anonymous):

of cos(x^2)?

OpenStudy (anonymous):

then your evaluate them then move on to the generate a formul

OpenStudy (anonymous):

part

OpenStudy (anonymous):

Yeah. But the first derivative is not -sin(x^2). You should use chain rule here. It should be: -2xsin(x^2)

OpenStudy (anonymous):

kk i didnt know that

OpenStudy (anonymous):

And for the second derivative, you should apply both the product rule and chain rule.

OpenStudy (anonymous):

I did it and found that all derivatives are zeros except those who are multiple of 4. I mean the 4th, 8th, 12th.. derivatives.

OpenStudy (anonymous):

kk sec....... im going to try to do it now and thanks for this your a life saver

OpenStudy (anonymous):

what did you get for the second derivative

OpenStudy (anonymous):

-4x^2cos(x^2)-2sin(x^2)

OpenStudy (anonymous):

the third derivative

OpenStudy (anonymous):

and you said the fourth derivative should be the one that isnt zero when evaluating it right?

OpenStudy (anonymous):

Yep. You should not bother yourself with finding these derivatives. I am pretty sure the question wants you to use the method I used.

OpenStudy (anonymous):

hey buddy also as i try to comprehend and understand this can you take a look at this Let f(x) = the cube root of x. Find the area of the surface generated by revolving the curve y = f(x) around the x-axis, for x ranging from 0 to 1.

OpenStudy (anonymous):

ok explain to me why it really doesnt matter to find the derivatives

OpenStudy (anonymous):

that is if you can

OpenStudy (anonymous):

You told me you know the Taylor series of cosx, then you just have to substitute for each x in that expansion of cosx by x^2. That's all.

OpenStudy (anonymous):

yes so the ans should be 1-1/2x^4+1/24x^8 then whats the nxt one

OpenStudy (anonymous):

-x^12/720

OpenStudy (anonymous):

ok now explain to me why not to worry about the derivative part because i must know this for tomorrow

OpenStudy (anonymous):

is there any tricks to getting the correct answere fast4r then going through each an evaluating it

OpenStudy (anonymous):

because after you find this i must know if it converges or diverges

OpenStudy (anonymous):

I mean you shouldn't worry about it in this particular problem, since you can find the Taylor series without doing any derivatives.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can you help me figure out if it now converges or diverges?

OpenStudy (anonymous):

You know how to use the ratio test?

OpenStudy (anonymous):

is there an easy way to know it so i know? or no

OpenStudy (anonymous):

because i dont understand exactly bc my professor did a poor job explaining it

OpenStudy (anonymous):

hmm I am not sure if there is a simpler way. All I know that you need to know the ratio test and the root test to find the radius and interval of convergence for power series such as Taylor series.

OpenStudy (anonymous):

is that*

OpenStudy (anonymous):

like ive got the definitions i front of me but explain why it converges or diverges

OpenStudy (anonymous):

ill talk to you on the other page this one i need for this page